# How Does the Smoothness Approximation Method Facilitate Generalization for Federated Adversarial Learning?

### v1

@@ Appendix F Generalization Analysis of SFAL
$=$ $\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\mathbb{E}_{\mathcal{S}}\left[\ell_{\rho}\left(\mathcal{A}(\mathcal{S});z_{i,j}\right)\right]=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\mathbb{E}_{\mathcal{S},z_{i^{\prime},j}^{\prime}}\left[\ell_{\rho}\left(\mathcal{A}\left(\mathcal{S}^{(i)}\right);z_{i^{\prime},j}^{\prime}\right)\right]$
@@
$\leq$ $\frac{(1+\alpha)\hat{m}}{\tilde{m}}\epsilon$
@@
$\leq$ $\frac{(1-\alpha)(m-\hat{m})}{\tilde{m}}\epsilon$
@@
If the Assumptions 1-4 hold and for any learning rate satisfying $\eta\leq\frac{1}{2\beta K}$, we can find the per-round recursion as
$\mathbb{E}[R(\theta_{t+1})]-\mathbb{E}[R(\theta_{t})]\leq$ $-\frac{K\eta}{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}+\frac{\eta\beta^{2}}{\tilde{m}}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\phi^{t}_{(\alpha,i)}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}+K\eta(\xi^{2}+\xi L)+\frac{2\beta K\eta^{2}\sigma^{2}}{\tilde{m}}$ (20)
@@
$\leq$ $\frac{2\beta K\eta^{2}\sigma^{2}}{\tilde{m}}+\dfrac{\beta mK\eta^{2}}{\tilde{m}^{2}}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}(\phi^{t}_{(\alpha,i)})^{2}\mathbb{E}\|\nabla R_{i}(\theta^{t+1}_{i,k})\|^{2}$
In the second inequality, we used (4) for the first and the second item. In the third inequality, we used Assumption 4 and $\frac{\sum_{i=1}^{m}(\phi^{t}_{(\alpha,i)})^{2}}{\tilde{m}}\leq 2$ for the first item.
@@
$+\frac{2\beta K\eta^{2}\sigma^{2}}{\tilde{m}}-\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\dfrac{\phi^{t}_{(\alpha,i)}\eta}{2\tilde{m}}\left(1-\frac{m\phi^{t}_{(\alpha,i)}}{2\tilde{m}}2\beta K\eta\right)\mathbb{E}\|\nabla R_{i}(\theta_{i,k}^{t+1})\|^{2}$
@@
Theorem 6. Given $\alpha\in[0,1),\hat{m}\leq\frac{m}{2},r_{\alpha}=1+\frac{\alpha}{1-\alpha}\frac{2\hat{m}}{m}$, we can obtain the following results in SFAL: 1. Under Assumption 1, let the step size be chosen as $\eta_{t}\leq\frac{\tilde{m}}{m\beta K(t+1)}$, the generalization bound $\varepsilon_{gen}$ with the surrogate smoothness approximation satisfies:
@@
2. Under Assumption 1, let $c\geq 0$ be a constant and the step size be chosen as $\eta_{t}\leq\frac{\tilde{m}\gamma}{4mK\sqrt{d}cL(t+1)}$. Under Assumption 1, the generalization bound $\varepsilon_{gen}$ with the randomized smoothness approximation satisfies:
@@
3. Without loss of generality, we assume $4K\eta_{t}C_{0}\geq 1$ and $s\geq 16\eta_{t}^{2}T^{2}K^{2}(b^{\prime}H_{K})^{2}(1+2\eta_{t}\zeta_{\theta})^{2}$,where $b^{\prime}=C_{x}^{2}B_{\varphi^{\prime\prime}}(C_{x}B_{\varphi^{\prime}}+\sqrt{2C_{0}}),H_{K}=2\sqrt{K\eta_{t}C_{0}}$. Let the step size be chosen as $\eta_{t}\leq\frac{\tilde{m}}{6mK\zeta_{\theta}}$. Under Assumptions 1-3, the generalization bound $\varepsilon_{gen}$ with the over-parameterized smoothness approximation satisfies:
@@
$\mathbb{E}\|\theta_{i,K}-\theta_{i,K}^{\prime}\|\leq e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+K\eta_{t}e^{\beta K\eta_{t}}(\xi+\frac{2}{n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|)$
@@
$\leq$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\beta K\eta_{t}}K\eta_{t}\xi+\dfrac{(1+\alpha)2K\eta_{t}e^{\beta K\eta_{t}}}{\tilde{m}n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|$
@@
$\leq$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\beta K\eta_{t}}K\eta_{t}\xi+\dfrac{2K\eta_{t}e^{\beta K\eta_{t}}}{\tilde{m}n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|$
@@
$=$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\beta K\eta_{t}}K\eta_{t}\xi+\dfrac{2K\eta_{t}e^{\beta K\eta_{t}}}{(1+\frac{\alpha}{1-\alpha}r)mn_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|$

### v2

@@ Appendix F Generalization Analyses of SFAL
$=$ $\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\mathbb{E}_{\mathcal{S}}\left[\ell_{\rho}\left(\mathcal{A}(\mathcal{S});z_{i,j}\right)\right]=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\mathbb{E}_{\mathcal{S},z_{i^{\prime},j}^{\prime}}\left[\ell_{\rho}\left(\mathcal{A}\left(\mathcal{S}^{(i)}\right);z_{i^{\prime},j}^{\prime}\right)\right].$
@@
$\mathbb{E}_{\mathcal{S}}\left[R_{i}(\mathcal{A}(\mathcal{S}))\right]=\frac{1}{n_{i}}\sum_{j=1}^{n_{i}}\mathbb{E}_{\mathcal{S},z_{i^{\prime},j}^{\prime}}\left[\ell_{\rho}\left(\mathcal{A}(\mathcal{S});z_{i^{\prime},j}^{\prime}\right)\right],$
@@
$\leq$ $\frac{(1+\alpha)\hat{m}}{\tilde{m}}\epsilon.$
@@
$\leq$ $\frac{(1-\alpha)(m-\hat{m})}{\tilde{m}}\epsilon.$
@@
If the Assumptions 1 and 4 hold and for any learning rate satisfying $\eta\leq\frac{1}{2\beta K}$, we can find the per-round recursion as
$\mathbb{E}[R(\theta_{t+1})]-\mathbb{E}[R(\theta_{t})]\leq$ $-\frac{K\eta}{2}\mathbb{E}\|\nabla R(\theta_{t})\|^{2}+\frac{\eta\beta^{2}}{\tilde{m}}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\phi^{t}_{(\alpha,i)}\mathbb{E}\|\theta^{t+1}_{i,k}-\theta_{t}\|^{2}+K\eta(\xi^{2}+\xi L)+\frac{2\beta K\eta^{2}\sigma^{2}}{\tilde{m}}.$
@@
$\leq$ $\frac{2\beta K\eta^{2}\sigma^{2}}{\tilde{m}}+\dfrac{\beta mK\eta^{2}}{\tilde{m}^{2}}\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}(\phi^{t}_{(\alpha,i)})^{2}\mathbb{E}\|\nabla R_{i}(\theta^{t+1}_{i,k})\|^{2}.$
In the second inequality, we used Lemma 4 for the first and the second item. In the third inequality, we used Assumption 4 and $\frac{\sum_{i=1}^{m}(\phi^{t}_{(\alpha,i)})^{2}}{\tilde{m}}\leq 2$ for the first item.
@@
$+\frac{2\beta K\eta^{2}\sigma^{2}}{\tilde{m}}-\sum\limits_{i=1}^{m}\sum\limits_{k=0}^{K-1}\dfrac{\phi^{t}_{(\alpha,i)}\eta}{2\tilde{m}}\left(1-\frac{m\phi^{t}_{(\alpha,i)}}{2\tilde{m}}2\beta K\eta\right)\mathbb{E}\|\nabla R_{i}(\theta_{i,k}^{t+1})\|^{2}.$
@@
Theorem 6. Given $\alpha\in[0,1),\hat{m}\leq\frac{m}{2},r_{\alpha}=1+\frac{\alpha}{1-\alpha}\frac{2\hat{m}}{m}$, we can obtain the following results in SFAL:
1. Under Assumption 1 and 4, let the step size be chosen as $\eta_{t}\leq\frac{\tilde{m}}{m\beta K(t+1)}$, the generalization bound $\varepsilon_{gen}$ with the surrogate smoothness approximation satisfies:
@@
2. Let $c\geq 0$ be a constant and the step size be chosen as $\eta_{t}\leq\frac{\tilde{m}\gamma}{4mK\sqrt{d}cL(t+1)}$. Under Assumption 1 and 4, the generalization bound $\varepsilon_{gen}$ with the randomized smoothness approximation satisfies:
@@
3. Without loss of generality, we assume $4K\eta_{t}C_{0}\geq 1$ and $s\geq 16\eta_{t}^{2}T^{2}K^{2}(b^{\prime}H_{K})^{2}(1+2\eta_{t}\zeta_{\theta})^{2}$,where $b^{\prime}=C_{x}^{2}B_{\varphi^{\prime\prime}}(C_{x}B_{\varphi^{\prime}}+\sqrt{2C_{0}}),H_{K}=2\sqrt{K\eta_{t}C_{0}}$. Let the step size be chosen as $\eta_{t}\leq\frac{\tilde{m}}{6mK\zeta_{\theta}}$. Under Assumptions 1-4, the generalization bound $\varepsilon_{gen}$ with the over-parameterized smoothness approximation satisfies:
@@
$\mathbb{E}\|\theta_{i,K}-\theta_{i,K}^{\prime}\|\leq e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+K\eta_{t}e^{\beta K\eta_{t}}(\xi+\frac{2}{n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|).$
@@
$\leq$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\beta K\eta_{t}}K\eta_{t}\xi+\dfrac{(1+\alpha)2K\eta_{t}e^{\beta K\eta_{t}}}{\tilde{m}n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|,$
@@
$\leq$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\beta K\eta_{t}}K\eta_{t}\xi+\dfrac{2K\eta_{t}e^{\beta K\eta_{t}}}{\tilde{m}n_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|.$
@@
$=$ $e^{\beta K\eta_{t}}\mathbb{E}\|\theta_{t}-\theta_{t}^{\prime}\|+e^{\beta K\eta_{t}}K\eta_{t}\xi+\dfrac{2K\eta_{t}e^{\beta K\eta_{t}}}{(1+\frac{\alpha}{1-\alpha}r)mn_{i}}\mathbb{E}\|g_{i}(\theta_{i,k})\|,$